Type or paste question here

Be sure to answer all parts.

Carry out these operations as if they were calculations of experimental results, and express each answer in the correct units and with the correct number of significant figures:

(a) 7.0895 m + 0.4 m + 2.39 m = ?m

(b) 3.67 g − 2.1721 g = ?g

(c) 4.07 cm × 3.0739 cm = ?cm2

(d) (2 ×103 g + 6.399 g ) / (0.065 cm3 − 0.029 cm3) = ?g/cm3
Answer #1

(a) 7.0895 + 0.4 + 2.39 = 9.9 m

The smallest number of decimal places out of any of number is 1. Hence, final answer has 1 decimal place.

(b) 3.67 - 2.1721 = 1.50 g ( 2 decimal places)

Smallest number of decimal places between two number is 2 . Final answer has 2 decimal place.

(C)

In multiplication , least number of significant figures in any number of the problem determines the significant figures in the final answer.

4.07 has 3 significant figures

3.0739 has 5 significant figures.

Hence, final answer has 3 significant figures.

4.07× 3.0739 = 12.5 cm2 ( 3 significant figures)

d)

( 2× 103 g ) = 200 ( 1 significant figure) .

200 + 6.399 = 206 g ( 0 decimal point)

(0.065 - 0.029) = 0.036 ( 3 decimal places)

(206/0.036) = 5700 g/cm3 ( 2 significant figure).